OpenStudy (anonymous):

Find the point on the terminal side of θ = -3 pi /4 that has an x coordinate of -1. Show your work for full credit

OpenStudy (anonymous):

@hartnn @jdoe0001

OpenStudy (dumbcow):

\[\cos(-3\pi/4) = \sin(-3pi/4) thus point is (-1,-1)